write the standard form of the equation of the circle that is tangent to x= -2 and has its center at (2, -4)
Answer by RAY100(1637)

the standard form for a circle is,,,(x-h)^2 +(y-k)^2 = r^2
.
with (h,k) as center of circle,,and r as radius
.
with center at (2, -4),,,,,h= 2,,,k= -4,,,and radius is the x distance from center to x=-2,,,|(-2-2)| = |(-4)| = 4
.
substituting, (x-2)^2 + (y-(-4) )^2 =4^2 =16
.
(x-2)^2 +(y+4)^2 = 16,,,,answer
.
.
a rough sketch shows how this is possible on the x-y coordinate system.
.
.